Hi! I have trouble with this word problem. Can you help me?
A rectangle and square have the same width, but the rectangles length is 5m longer than the squares. Their area is 133m(squared). Find the dimensions.
I have tried
rectange=l+5
square= l
then I did l+l+5=133
-5=-5
2l=128
i divided by 2 and got
l=64m
I dont think that is right. Can you help me?
If you could help me that would be great, but if not I'll try to work it out differently.
Thanks,
Melissa

Answer by fractalier(6550)

Here's how you do it.
Let w be the width of both.
The square, then, has a length of w also.
The rectangle has a length of w + 5.
Now if we add the areas, we get a total of 133, so here goes...
w^2 + w(w + 5) = 133
2w^2 + 5w - 133 = 0
(2w + 19)(w - 7) = 0
w = -19/2 or w = 7
but distance must be positive so w = 7.
The length of the rectangle is 12.